# FUSION: Forecast-Embedded Agent Scheduling with Service Incentive Optimization over Distributed Air-Ground Edge Networks

### v1

@@ IV-C1 Design of Proactive Spatio-Temporal Demand Prediction via Liquid Neural Networks (Pro-LNN)
Unlike discrete-time RNNs or Transformers, LNNs [23] rely on exponential-style state updates with neuron-wise time constants, making them naturally suited for irregular sampling, non-stationary patterns, and lightweight edge deployment. In our setting, the service demand of each ES exhibits strong spatio–temporal variability with occasional missing or uneven observations. Hence, the predictor must remain stable under irregular effective time gaps (denoted by $\Delta\nu$) while capturing both short-lived bursts and long-term trends, where $\Delta\nu$ represents the time interval between two consecutive trading rounds. To avoid confusion with the fine-grained time-slot index used in the online stage, we use the Greek letter $\nu$ to index trading rounds throughout this subsection.
Let $N_{i}^{\mathrm{Dem},(\nu)}$ denote the task demand of ES $s_{i}$ at the $\nu$-th trading round, and $\mathbf{z}_{i}^{(\nu)}=[N_{i}^{\mathrm{Dem},(\nu)},\mathbf{x}_{i}^{(\nu)}]$ collect the demand and optional exogenous features. Given a length-$L$ history $\{\mathbf{z}_{i}^{(\nu-L+1)},\ldots,\mathbf{z}_{i}^{(\nu)}\}$, Pro-LNN encodes this sequence via a stable liquid cell with neuron-wise time constants, integrates the observations into a hidden state $\mathbf{h}_{i}^{(\nu)}$, and then applies a shallow readout network to produce an $H$-step forecast $\{\hat{N}_{i}^{\mathrm{Dem},(\nu+1)},\ldots,\hat{N}_{i}^{\mathrm{Dem},(\nu+H)}\}$. The model is trained with an $H$-step MSE loss, optionally augmented by a multi-step consistency term that rolls the predictor forward in closed loop and penalizes accumulated errors at future trading rounds. These forecasts are computed offline before each trading round and then fed into the eACO-VRP module for ES recommendation and AgP route pre-planning, as well as the auction module for ES–AgP contract design. The detailed update equations and training pseudocode of Pro-LNN are deferred to Appx. B (Algorithm 4) due to space limitations.
$p^{\kappa}_{\mathrm{n}_{r}\rightarrow\mathrm{n}_{s}}=\begin{cases}\displaystyle\frac{\varphi^{\varepsilon_{1}}_{r,s}\,\eta^{\varepsilon_{2}}_{r,s}\,\big(1/\mathrm{width}_{s}\big)^{\varepsilon_{3}}\,\big(1/\mathrm{wait}_{s}\big)^{\varepsilon_{4}}}{\sum\limits_{\mathrm{n}_{q}\in J_{\kappa}(\mathrm{n}_{r})}\varphi^{\varepsilon_{1}}_{r,q}\,\eta^{\varepsilon_{2}}_{r,q}\,\big(1/\mathrm{width}_{q}\big)^{\varepsilon_{3}}\,\big(1/\mathrm{wait}_{q}\big)^{\varepsilon_{4}}},&\text{if }\mathrm{n}_{s}\in J_{\kappa}(\mathrm{n}_{r}),\\[12.0pt] 0,&\text{otherwise.}\end{cases}$ (6)
@@ IV-C2 Enhanced Ant Colony Optimization for Vehicle Route Planning (eACO-VRP)
We denote $\mathbb{K}$ as the set of ants. Each ant $\mathbbm{k}\in\mathbb{K}$ simulates the trajectory of $v_{k}$, starting from $\mathrm{n}_{0}$. At each step, it evaluates: the current timeslot, ES time windows, vertex locations, the travel time to each candidate ES, and the utility upon arrival; it then selects a feasible next vertex to visit (i.e., whether the arrival time from the current vertex plus travel time remains before the end of the candidate vertex’s time window). This process continues until no visitable vertex remains. The state-transition rule, i.e., the probability that ant $\mathbbm{k}$ moves from $\mathrm{n}_{r}$ to $\mathrm{n}_{s}$ ($r\neq s\neq 0$), is given by (6), with positive weights $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\varepsilon_{4}$. We denote by $\varphi_{r,s}$ the pheromone level on edge $(\mathrm{n}_{r},\mathrm{n}_{s})$ (see (7)); $\eta_{r,s}$ is the distance between $\mathrm{n}_{r}$ and $\mathrm{n}_{s}$; the time-window width at vertex $\mathrm{n}_{s}$ is $\mathrm{width}_{s}=t^{\mathrm{e}}_{s}-t^{\mathrm{b}}_{s}$. If the ant arrives early (e.g., before the ES demand window starts), the waiting time $\mathrm{wait}_{s}$ is defined as the arrival time minus the starting time of the ES’s demand window. Let $J_{\mathbbm{k}}(r)$ be the set of feasible ES vertices that ant $\mathbbm{k}$ can still complete when standing at node $\mathrm{n}_{r}$. An ES $\mathrm{n}_{s}\in J_{\mathbbm{k}}(r)$ must satisfy two feasibility conditions: (i) the current time plus the travel time from $\mathrm{n}_{r}$ to $\mathrm{n}_{s}$ does not exceed the starting time of $s$’s demand window, and (ii) the current UAV load of AgP $v_{k}$ is sufficient to satisfy the demand of $\mathrm{n}_{s}$.
Once a route is found by the colony, pheromones on edges are updated by
where $0<\theta<1$ is the evaporation coefficient, and the deposited pheromone is
$\Delta\varphi_{r,s}=\left\{\begin{aligned} &\sum_{(\mathrm{n}_{r},\mathrm{n}_{s})\in\bm{M}^{\mathrm{max}}}\,\triangle{U^{\mathrm{V}}_{r,s}},\quad&&\text{if }(\mathrm{n}_{r},\mathrm{n}_{s})\in\bm{M}^{\mathrm{max}},\\ &0,&&\text{otherwise},\end{aligned}\right$. (8)
where $\triangle{U^{\mathrm{V}}_{r,s}}$ denotes the utility increment for $v_{k}$ when moving from $\mathrm{n}_{r}$ to serve the ES at $\mathrm{n}_{s}$. The path with the largest utility among all ants is $\bm{M}^{\mathrm{max}}=\Big\{\bm{M}^{\mathbbm{k}}\ \Big|\ \max~\sum_{(r,s)\in\bm{M}^{\mathbbm{k}}}\triangle{U^{\mathrm{V}}_{r,s}},\ \forall\,\mathbbm{k}\in\mathbb{K}\Big\}$. Alg. 1 details the pseudo-code of eACO-VRP:
[…]
2
[…]
4
5while $iter\leq iter_{\mathrm{max}}$ do
6  for each ant $\kappa\in\mathbb{K}$ do
7     $\bm{M}^{\kappa}\!\leftarrow\!\varnothing$;  $\mathrm{n}_{\text{cur}}\!\leftarrow\!\mathrm{n}_{0}$
8     while end condition not met do
9        Compute: feasible set $J_{\kappa}(\mathrm{n}_{\text{cur}})$ and transition probabilities $p^{\kappa}_{\mathrm{n}_{\text{cur}}\rightarrow\mathrm{n}_{s}}$ (via (6))
10        if $J_{\kappa}(\mathrm{n}_{\text{cur}})\neq\varnothing$ then
11           sample next node $\mathrm{n}_{\text{next}}\sim p^{\kappa}_{\mathrm{n}_{\text{cur}}\rightarrow\mathrm{n}_{s}}$ over $\mathrm{n}_{s}\!\in\!J_{\kappa}(\mathrm{n}_{\text{cur}})$
12           $\bm{M}^{\kappa}\!\leftarrow\!\bm{M}^{\kappa}\cup\{\text{edge}(\mathrm{n}_{\text{cur}},\mathrm{n}_{\text{next}})\}$
13           $\mathrm{n}_{\text{cur}}\!\leftarrow\!\mathrm{n}_{\text{next}}$
14       else
15          break
16
17
18
19 $\bm{M}^{\mathrm{max}}\!\leftarrow\!\arg\max_{\bm{M}^{\kappa}:\kappa\in\mathbb{K}}\;\sum_{(\mathrm{n}_{r},\mathrm{n}_{s})\in\bm{M}^{\kappa}}\triangle\overline{U^{\mathrm{W}}_{r,s}}$
[…]
21    Pheromone evaporation and deposit (7) and (8)
22
23 $iter\!\leftarrow\!iter+1$
24
Algorithm 1 Proposed eACO-VRP
[…]
@@ 23 $iter\!\leftarrow\!iter+1$
Step 2. Path exploration. At each iteration, an ant $\mathbbm{k}$ located at its current vertex $\mathrm{n}_{r}$ jointly evaluates the current time, the travel time to each unvisited vertex, the service time at that vertex, and the corresponding time-window constraint to determine feasibility. This screening step yields the feasible candidate set $J_{\mathbbm{k}}(r)$ and the associated transition probabilities $pr^{\mathbbm{k}}_{r,s}$ (line 7, Alg. 1). The ant then stochastically selects its next vertex from $J_{\mathbbm{k}}(r)$ according to (6). The construction of a path for ant $\mathbbm{k}$ terminates once no executable vertex remains (line 13, Alg. 1).
@@
Step 3. Pheromone update. For each constructed path $\bm{M}^{\mathbbm{k}}$, we compute the cumulative utility $\sum_{(r,s)\in\bm{M}^{\mathbbm{k}}}\triangle{U^{\mathrm{V}}_{r,s}}$. Among all ants, the path with the largest utility is selected as $\bm{M}^{\mathrm{max}}$, and the pheromone trails on its edges are reinforced (lines 14-16, Alg. 1) according to the update rules in (7) and (8), while pheromones on other edges evaporate.
@@
Step 4. Output. After all iterations, the route $\bm{M}^{\mathrm{max}}$ with the highest utility is returned as the pre-planned path for AgP $v_{k}$, and the ES tasks located along this route are recommended as its preferred service targets.
1Input: ESs $\bm{\mathcal{S}}$ with demands $\{N_{i}^{\mathrm{Dem}}\}$ and bid matrix $\{\mathrm{bid}_{i,k}\}$; AgPs $\bm{\mathcal{V}}$ with asks $\{\mathrm{ask}_{k}\}$ and capacities $\{C_{k}^{\mathrm{V}}\}$; auxiliary info $\{\mathbbm{a}_{i},\mathbbm{d}_{k},\mathbbm{r}_{k}\}$; parameter $\mu$.
2[0.1em]
3Initialization: Compute per-ES reference bids $\overline{\mathrm{bid}}_{i}\leftarrow\mathrm{mean}_{k}(\mathrm{bid}_{i,k})$;
4 Utility matrix $U_{i,k}\leftarrow\mathrm{bid}_{i,k}-\mathrm{ask}_{k}$;
5 Set $\bm{X}^{*}\!\leftarrow\!\bm{0}$, $\bm{p}^{\mathrm{b}*}\!\leftarrow\!\bm{0}$, $\bm{r}^{\mathrm{s}*}\!\leftarrow\!\bm{0}$.
6# Phase 1: List generation & pivotal-rank search
7 Sort ES indices into $\mathcal{L}^{\mathrm{ES}}$ by $\overline{\mathrm{bid}}_{i}$ (desc),
8    and AgP indices into $\mathcal{L}^{\mathrm{AgP}}$ by $\mathrm{ask}_{k}$ (asc); relabel $(i,k)$ accordingly.
9 Find a pair of pivotal ranks $(I^{\ast},K^{\ast})$ such that $\overline{\mathrm{bid}}_{I^{\ast}}\geq\mathrm{ask}_{K^{\ast}},\quad\sum_{k=1}^{K^{\ast}}C_{k}^{\mathrm{V}}\;\geq\;\sum_{i=1}^{I^{\ast}}N_{i}^{\mathrm{Dem}}$, and $I^{\ast}$ is maximized (breaking ties by larger capacity slack $\sum_{k=1}^{K^{\ast}}C_{k}^{\mathrm{V}}-\sum_{i=1}^{I^{\ast}}N_{i}^{\mathrm{Dem}}$).
10# Phase 2: ES–AgP matching
11 Initialize the set of unmatched ESs: $\mathcal{U}^{\mathrm{free}}\leftarrow\{i\mid 1\leq i\leq I^{\ast}\}$.
12 for $k$ in top-$K^{\ast}$ AgPs (according to $\mathcal{L}^{\mathrm{AgP}}$) do
13  if $C_{k}^{\mathrm{V}}\leq 0$ or $\mathcal{U}^{\mathrm{free}}=\emptyset$ then
14     continue
15 Build candidate ES set $\mathcal{C}_{k}\leftarrow\big\{i\in\mathcal{U}^{\mathrm{free}}\,\big|\,U_{i,k}>0\ \wedge\ \mathrm{bid}_{i,k}\geq\overline{\mathrm{bid}}_{i}\big\}$;
16 if $\mathcal{C}_{k}=\emptyset$ then
17     continue
18 $\mathcal{K}_{k}^{\star}\leftarrow$ Alg. 1 (eACO-VRP) with input $(\mathcal{C}_{k},\{N_{i}^{\mathrm{Dem}}\}_{i\in\mathcal{C}_{k}},C_{k}^{\mathrm{V}})$;
19  for $i\in\mathcal{K}_{k}^{\star}$ do
20     Set $x^{\mathrm{Off}}_{i,k}\leftarrow 1$;
21     Remove $i$ from $\mathcal{U}^{\mathrm{free}}$: $\mathcal{U}^{\mathrm{free}}\leftarrow\mathcal{U}^{\mathrm{free}}\setminus\{i\}$;
22 Update remaining capacity: $C_{k}^{\mathrm{V}}\leftarrow C_{k}^{\mathrm{V}}-\sum_{i\in\mathcal{K}_{k}^{\star}}N_{i}^{\mathrm{Dem}}$.
23Denote the resulting matching by $\bm{X}^{*}$.
24# Phase 3: Pricing via monotone binary search (contract determination)
25 (ES-side payments)
26 for each matched ES $i$ with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$ do
27  Set $high\leftarrow\overline{\mathrm{bid}}_{i}$, $low\leftarrow\overline{\mathrm{bid}}_{I^{\ast}+1}$ (or a safe reserve);
28  Let $b_{i}^{\mathrm{eff}}\leftarrow\overline{\mathrm{bid}}_{i}$ be ES $i$’s effective bid parameter.
29  while $low<high$ do
30     Temporarily set $b_{i}^{\mathrm{eff}}\leftarrow(low+high)/2$;
31     Rerun Phase 2 on the top-$I^{\ast}$ ESs and top-$K^{\ast}$ AgPs (using $b_{i}^{\mathrm{eff}}$ for ES ii) to obtain $\bm{X}^{\mathrm{Temp}}$;
32     if $\sum_{k}x^{\mathrm{Temp}}_{i,k}=1$ then $high\leftarrow b_{i}^{\mathrm{eff}}$;
33     else $low\leftarrow b_{i}^{\mathrm{eff}}$;
34
35 Set $p_{i}^{\mathrm{b}*}\leftarrow high$.
36
37(AgP-side rewards)
38 for each matched AgP $k$ with $\sum_{i}x^{\mathrm{Off}}_{i,k}\geq 1$ do
40  while $low<high$ do
41     Temporarily set $\mathrm{ask}_{k}^{\mathrm{eff}}\leftarrow(low+high)/2$;
42     Rerun Phase 2 with $\mathrm{ask}_{k}^{\mathrm{eff}}$ to obtain $\bm{X}^{\mathrm{Temp}}$;
43     if $\sum_{i}x^{\mathrm{Temp}}_{i,k}\geq 1$ then $low\leftarrow\mathrm{ask}_{k}^{\mathrm{eff}}$;
44     else $high\leftarrow\mathrm{ask}_{k}^{\mathrm{eff}}$;
45
46 Set $r_{k}^{\mathrm{s}*}\leftarrow low$.
47
Output: Matching $\bm{X}^{*}=\{x^{\mathrm{Off}}_{i,k}\}$, payments $\bm{p}^{\mathrm{b}*}$, rewards $\bm{r}^{\mathrm{s}*}$.
Algorithm 2 Proposed Off-AIC2
@@ IV-C3 Offline Auction-based Incentive-Compatible Contracting (Off-AIC2)
Following the pre-planned routes, we then introduce Off-AIC2 (Alg. 2), which realizes efficient and incentive-aware offline contracts between ESs and AgPs, inspired by double auction mechanisms in distributed markets [22].
@@
Step 1. List generation and pivotal-index search. Off-AIC2 starts as a sealed-bid, private, and collusion-free mechanism, where all ES buyers $s_{i}$ and AgP sellers $s_{k}$ simultaneously submit their bids and asks to the auctioneer. We first sort participants to align the bid–ask hierarchy: ESs are ordered into a list $\mathcal{L}^{\mathrm{ES}}$ by descending average bid $\overline{\mathrm{bid}}_{i}$, and AgPs are ordered into a list $\mathcal{L}^{\mathrm{AgP}}$ by ascending ask price $\mathrm{ask}_{k}$. Without loss of generality, we relabel indices so that these sorted orders are reflected directly by $i$ and $k$. Then, the algorithm searches for a pair of pivotal ranks $(I^{\ast},K^{\ast})$ such that the cumulative capacity of the top-$K^{\ast}$ AgPs can cover the aggregate demand of the top-$I^{\ast}$ ESs under a feasible price relationship as $\overline{\mathrm{bid}}_{I^{\ast}}\geq\mathrm{ask}_{K^{\ast}},\sum_{k=1}^{K^{\ast}}C_{k}^{\mathrm{V}}\geq\sum_{i=1}^{I^{\ast}}N_{i}^{\mathrm{Dem}}$. Intuitively, $(I^{\ast},K^{\ast})$ identifies a largest feasible trading region where high-bidding ESs are matched with low-asking AgPs, and the selected AgPs have sufficient aggregate capacity to serve the selected ESs, thus laying the foundation for near-clearing and budget-balanced trading under current market conditions.

### v2

@@ IV-C2 Enhanced Ant Colony Optimization for Vehicle Route Planning (eACO-VRP)
$pr^{\mathbbm{k}}_{r,s}=\begin{cases}\displaystyle\frac{\varphi^{\varepsilon_{1}}_{r,s}\,\eta^{\varepsilon_{2}}_{r,s}\,\big(1/\mathrm{width}_{s}\big)^{\varepsilon_{3}}}{\sum\limits_{q\in J^{\mathbbm{k}}(\mathrm{n}_{r})}\varphi^{\varepsilon_{1}}_{r,q}\,\eta^{\varepsilon_{2}}_{r,q}\,\big(1/\mathrm{width}_{q}\big)^{\varepsilon_{3}}}&\text{if }\mathrm{n}_{s}\in J^{\mathbbm{k}}(\mathrm{n}_{r}),\\[12.0pt] 0&\text{otherwise,}\end{cases}$
with positive weights $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}$. In (6), $\varphi_{r,s}$ denotes the pheromone level on edge $(\mathrm{n}_{r},\mathrm{n}_{s})$ (see (7)); $\eta_{r,s}$ is the distance between $\mathrm{n}_{r}$ and $\mathrm{n}_{s}$. If the ant arrives early (i.e., before the time-slot when ES $s_{i}$ begins to request extra resources), it can pre-position the AP’s UAV at this ES; in this case, the remaining time $\mathrm{width}_{s}$ is defined as the difference between the arrival time and the time-slot when ES $s_{i}$ begins to request extra resources. Conversely, if the ant reaches $\mathrm{n}_{s}$ after the time-slot when ES $s_{i}$ begins to request extra resources, this ES is regarded as infeasible and eliminated in the route construction. In (6), $J^{\mathbbm{k}}(\mathrm{n}_{r})$ denotes the set of feasible ES vertices that ant $\mathbbm{k}$ can reach when standing at node $\mathrm{n}_{r}$: an ES $\mathrm{n}_{s}\in J^{\mathbbm{k}}(\mathrm{n}_{r})$ if it satisfies two feasibility conditions: (i) the current time plus the travel time from $\mathrm{n}_{r}$ to $\mathrm{n}_{s}$ should not exceed the time-slot when ES $s_{i}$ begins to request extra resources, and (ii) there must still be remaining UAVs at AP $v_{k}$ to deploy. Once a route is found by ants, pheromones on edges are updated as follows:
where $0<\theta<1$ is the evaporation coefficient, and
$\Delta\varphi_{r,s}=\left\{\begin{aligned} &\sum_{(\mathrm{n}_{r},\mathrm{n}_{s})\in\bm{M}^{\mathrm{max}}}\,\Delta{U^{\mathrm{A}}_{r,s}},\quad&&\text{if }(\mathrm{n}_{r},\mathrm{n}_{s})\in\bm{M}^{\mathrm{max}},\\ &0,&&\text{otherwise},\end{aligned}\right$. (8)
captures the deposited pheromone, where $\triangle{U^{\mathrm{A}}_{r,s}}$ denotes the utility increment for $v_{k}$ when moving from $\mathrm{n}_{r}$ to serve the ES at $\mathrm{n}_{s}$, and $\bm{M}^{\mathrm{max}}=\Big\{\bm{M}^{\mathbbm{k}}\ \Big|\ \max~\sum_{(r,s)\in\bm{M}^{\mathbbm{k}}}\triangle{U^{\mathrm{A}}_{r,s}},\ \forall\,\mathbbm{k}\in\mathbb{K}\Big\}$ denotes the path with the largest utility among ants. For brevity, the pseudo-code of eACO-VRP is provided in Appx. B-C, while its four main steps are summarized below:
@@
Step 2. Path exploration: At each iteration, an ant $\mathbbm{k}$ located at its current vertex $\mathrm{n}_{r}$ jointly evaluates the current time-slot, the travel time to each vertex, the service time at that vertex, and the time-slot when ES $s_{i}$ begins to request extra resources. This screening step yields the feasible candidate set $J^{\mathbbm{k}}(\mathrm{n}_{r})$ and the associated transition probabilities $pr^{\mathbbm{k}}_{r,s}$ as discussed above. The ant then probabilistically selects its next vertex using (6).
@@
Step 3. Pheromone update: For each constructed path $\bm{M}^{\mathbbm{k}}$, we compute the cumulative utility $\sum_{(r,s)\in\bm{M}^{\mathbbm{k}}}\triangle{U^{\mathrm{A}}_{r,s}}$. Among all ants, the path with the largest utility is selected as $\bm{M}^{\mathrm{max}}$, and the pheromone trails on its edges are reinforced according to (7) and (8), while pheromones on other edges evaporate.
@@
Step 4. Output: After finishing all the iterations, the route $\bm{M}^{\mathrm{max}}$ with the highest utility is returned as the pre-planned path for AP $v_{k}$, and the ES tasks located along this route are recommended as its preferred service targets.
@@ IV-C3 Offline Auction-based Incentive-Compatible Contracting (Off-AIC2)
Building on the pre-planned AP routes obtained above, we now introduce Off-AIC2, an offline double-auction mechanism that establishes incentive-aware service contracts between ESs and APs. The goal is to decide ES–AP associations, how much service capacity should be reserved along each AP route, and how payments/rewards should be set so that participants (i.e., ESs and APs) have limited incentive to misreport their private values.
In Off-AIC2, on the buyer side, each ES $s_{i}$ predicts its total service demand $N_{i}^{\mathrm{Dem}}$ (e.g., the number of service units or tasks that need to be supported in the next trading round) and has a private per-unit valuation $v_{i}$ for receiving edge service. It then participates in Off-AIC2 by submitting a per-unit bid $\overline{\mathrm{bid}}_{i}$ together with its demand $N_{i}^{\mathrm{Dem}}$, where $\overline{\mathrm{bid}}_{i}$ represents the maximum payment it is willing to make per unit of service. On the seller side, each AP $v_{k}$, based on its maximum service capacity $C_{k}^{\mathrm{A}}$ (e.g., the total number of service units it can provide) and its service cost, reports a per-unit ask price $\mathrm{ask}_{k}$ and its capacity $C_{k}^{\mathrm{A}}$, where $\mathrm{ask}_{k}$ is the minimum acceptable payment per unit of service. Subsequently, for a potential ES–AP pair $(s_{i},v_{k})$, we define a pairwise utility $U_{i,k}^{\mathrm{Off}}=\mathrm{bid}_{i,k}-\mathrm{ask}_{k}$ ($U_{i,k}^{\mathrm{Off}}>0$ indicates a socially beneficial match). Off-AIC2 then uses the reported bids, asks, demands, and capacities to compute pairwise utilities and select a set of winning ESs and APs, determine their matching/association pattern, and compute payments and rewards. For brevity, the complete pseudo-code of Off-AIC2 is provided in Appx. B-D, while its main three steps are summarized below:
@@
Step 1. List generation and pivotal-index search: Off-AIC2 operates as a one-shot sealed-bid double auction: all ES buyers $s_{i}$ and AP sellers $v_{k}$ simultaneously submit their bids, asks, and resource quantity information to an auctioneer without observing each others’ submissions. We first sort ESs into a list $\mathcal{L}^{\mathrm{ES}}$ in descending order of $\overline{\mathrm{bid}}_{i}$, and APs into a list $\mathcal{L}^{\mathrm{AP}}$ in ascending order of $\mathrm{ask}_{k}$. Without loss of generality, we relabel indices so that these sorted orders are directly reflected by $i$ and $k$. The algorithm then searches for a pair of pivotal ranks $(I^{\ast},K^{\ast})$ such that $\overline{\mathrm{bid}}_{I^{\ast}}\geq\mathrm{ask}_{K^{\ast}},\sum_{k=1}^{K^{\ast}}C_{k}^{\mathrm{A}}\geq\sum_{i=1}^{I^{\ast}}N_{i}^{\mathrm{Dem}}$.
In words, $(I^{\ast},K^{\ast})$ is the largest prefix of high-bidding ESs and low-asking APs whose aggregate demand and supply can be satisfied at a non-loss-making price level.
Step 2. ES–AP matching and route planning via eACO-VRP: Given the pivotal ranks $(I^{\ast},K^{\ast})$, the top-$I^{\ast}$ ESs and top-$K^{\ast}$ APs participate in the matching stage. We initialize the set of currently unmatched ESs as $\mathcal{U}^{\mathrm{free}}=\{1,\dots,I^{\ast}\}$ and iterate over AP sellers $v_{k}$ following the order in $\mathcal{L}^{\mathrm{AP}}$. For each AP $v_{k}$ with remaining capacity $C_{k}^{\mathrm{A}}>0$ and $\mathcal{U}^{\mathrm{free}}\neq\emptyset$, we construct a candidate ES set $\mathcal{C}_{k}=\big\{\,s_{i}\in\mathcal{U}^{\mathrm{free}}\,\big|\,U_{i,k}^{\mathrm{Off}}>0,\ \overline{\mathrm{bid}}_{i}\geq\mathrm{ask}_{k}\big\}$, where $U_{i,k}^{\mathrm{Off}}>0$ ensures that serving $s_{i}$ along some route of $v_{k}$ improves the system-level SW, and $\overline{\mathrm{bid}}_{i}\geq\mathrm{ask}_{k}$ guarantees that the ES’s per-unit value is larger than the AP’s ask price.